OpenStudy (anonymous):

i need some help Identifying the 12th term of a geometric sequence where a1 = 8 and a6 = –8,192.

OpenStudy (anonymous):

\[ a_1= 8\\ a_2= 8 r\\ a_3= 8 r^2\\ a_4= 8 r^3\\ a_5 = 8 r^4\\ a_6= 8 r^5=8,192.\\ r=-4 \]

OpenStudy (anonymous):

\[ a_{12}= 8 (-4)^{11}= -33554432 \]